A cake retailer increases the price of its cakes from £2.00 to £2.60 per cake and the quantity demanded decreases from 60 cakes per day to 45 cakes per day.

The price elasticity of demand for the retailer’s cake is:

A​-0.25

B​-0.3

C​-0.83

D​-1.2


1
Expert's answer
2022-03-13T19:01:29-0400

Price elasticity of demand = percentage change in quantity/percentage change in price

percentage change in quantity = "\u200b\n \\frac{60-45}{{105}\/2}"


="\u200b\n \\frac{15}{52.5}*100"

=28.57%

percentage change in price = "\u200b" "\u200b\n \\frac{2.60-2.00}{4.6\/2}"

​ = "\u200b\n \\frac{0.6}{2.3}""*100"

=26.09%

price elasticity of demand = "\u200b\n \\frac{28.57}{26.09}" = -1.095

the answer is approximately D.

"\u200b"